OpenStudy (anonymous):

a sled is pushed along an ice covered lake. it has some initial velocity before coming to a rest in 15m. it took 23 seconds before the sled and rider come to a rest. if the rider and sled have a combined mass of 52.5 kg, what is the magnitude and direction of the stopping force? what do we call the stopping force?

OpenStudy (anonymous):

F=m.a =m(v-u)/t (since v=o, a=-u/t) v-u=d/t -u=15/23 u=-0.65 m/s F=m.-u/t =52.5*0.65 =1.47 N Conventionally, for the stopping force, F is taken to be negative. So, the stopping force is -1.47 N and this force is the frictional force and its direction is opposite to that of motion.